OpenStudy (anonymous):

A landscaper is designing a flower garden in the shape of a trapezoid.She wants the shorter base to be 3 yards greater than the height and the longer base to be 7 yards greater than the height. She wants the area to be 295 square yards.The situation is modeled by the equation h^2+5h=295.Use the quadratic formula to find the height that will give the desired area.Round to the nearest hundredth of a yard. A)17.36 yards B)600 yards C)14.86 yards D)29.71 yards can someone help me?

OpenStudy (anonymous):

Well eh.. Area of trapezoid is the (average of bases) * (height) So, if we mark the height as 'x' we can say: \[ \frac{(x + 3) + (x + 7)}{2} \cdot x = 295 \\ \frac{2x + 10}{2} \cdot x = 295 \\ (x + 5)x = 295 \\ x^2 + 5x - 295 = 0 \\ x_{1,2} = \frac{ -5 \pm \sqrt{ 25 + 4 \cdot 295 } }{2} \\ x_{1,2} = \frac{ -5 \pm \sqrt{ 1205 } }{2} \\ x_{1,2} = \frac{ -5 \pm 34.71 }{2} \\ x_{1} = \frac{ -5 + 34.71 }{2} = \frac{ 29.71 }{2} = 14.86\\ x_{2} = \frac{ -5 - 34.71 }{2} = - \frac{ 39.71 }{2} \] But since x2 is negative.. and height can't be negative, we're left with x1, which is roughly 14.86

OpenStudy (anonymous):

Thank you @pitamar